myininaya (myininaya):

here's an easy one no cheating don't go to wolframalpaha \int_{}^{} \frac{e^{x}}{1+e^{2x}} dx here's an easy one no cheating don't go to wolframalpaha \int_{}^{} \frac{e^{x}}{1+e^{2x}} dx @Mathematics

myininaya (myininaya):

\[\int\limits_{}^{} \frac{e^{x}}{1+e^{2x}} dx \]

OpenStudy (anonymous):

substitution!

myininaya (myininaya):

yes!

OpenStudy (anonymous):

let e^x = u

myininaya (myininaya):

then we have arctan(u)+C =arctan(e^x)+C gj :)

myininaya (myininaya):

see easy :)

OpenStudy (karatechopper):

how? im in 7th!!!